Robert is now twice as old as Marie. Six years ago the sum of their ages was thirty-nine. How old are they each now?
Answer by ewatrrr(24785)

 
Hi,
Robert is now twice as old as Marie
Let x and 2x represent Marie's and Robert's ages NOW
Question states*** 6 yrs ago...
(x-6) + (2x-6) = 39
Solving for x
3x = 51
x = 17, Marie's age now. Robert is 34 (2*17)
CHECKING our Answer***
11 + 28 = 39
One side of a triangle is three times the shortest side.
The third side is 6 inches more than the shortest side.
Let x, 3x and (x+6) represent the three sides
Question states*** The perimeter of the triangle is 86 inches
x + 3x + (x+6) = 86in
5x = 80in
x = 16in. Sides are 16in, 48in, 22in (added up = 86in :)